OpenStudy (anonymous):

Find the antiderivative of sin(2t+5) and e^5x+2x-3

zepdrix (zepdrix):

\[\Large \int\limits \sin(2t+5)\;dt\]Mmmmm you can do a u-sub if you want, but guessing is easier imo. What's the anti-derivative of sine?

OpenStudy (anonymous):

-cos

zepdrix (zepdrix):

\[\Large \int\limits\limits \sin(2t+5)\;dt\quad\approx\quad -\cos(2t+5)+C\]Ok so there is our guess. If we take the derivative, do we end up with sin(2t+5)? How much are we off by?

zepdrix (zepdrix):

Where you at sammmm? :P Don't like this method?

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

you end up with -1/2cos(2t+5)

zepdrix (zepdrix):

Yes you do :3

OpenStudy (anonymous):

why doesn't anything change in the parentheses?

OpenStudy (anonymous):

using substitution you get u = 2t + 5 du = 2dt which means dt = du/2 hence your equation becomes \[\int\limits_{}^{}\frac{ \sin(u)}{ 2 }du\] now integrating with respect to u gives \[\frac{ -\cos(u) }{ 2 }\] and finally you get \[\frac{ -\cos(2t+5) }{ 2 }\]

zepdrix (zepdrix):

Why doesn't the inner function change? Hmm taking the derivative should give you an indication as to why D: Nothing happens to the inner function when we take a derivative right? It's the same idea with integration :d

OpenStudy (anonymous):

so is -cos(2t+5)/2+C the same as -1/2cos(2t+5)+C

zepdrix (zepdrix):

The same? :o no... It's a different method for solving these types of problems. I find u-subs to be annoying.. So I use a method of advanced guessing, and then you just adjust your solution at the end. But do whatever works for you :)

zepdrix (zepdrix):

Oh sorry, i didn't see the 2 you wrote :) yes those are the same!

OpenStudy (anonymous):

what if I were to find the derivative instead of the antiderivative?